A student opens a book to two facing pages. The product of the page numbers is 2550. What are thepage numbers?
Answer by mananth(16946)

let the left side page number be x
so the next page number will be x+1
x(x+1)= 2550
x^2+x=2550
x^2+x-2550=0
x^2+51x-50x-2550=0
x(x+51)-50(x+51)=0
(x+51)(x-50)=0
x= 50 ignore the negative value 51
the pages are 50 & 51